let n€N consider the set of nxn symmetric matrices over R with the usual addition and multiplication by a scalar. show that this set with the given operations is a vector subspace of Mnn

1
Expert's answer
2021-11-03T18:33:09-0400



We have

"V=M_n(\R)"


is the vector space over "\R" .


Now,

denote "A_{ij}" is the matrix whose "(i,j)^{th}" entry is "1" if "i=j"

and ;if


"i\ne j,\,\forall i,j\in \{1,\dots,n\}"

Also denote the collection of all such matrix by "X" ,

thus;


"W=\text{span}(X)" is the smallest subspace


such that "I_{n\times n}\in W" .


Define


"M_n^s(\R):=\{A\in M_n(\R):A=A^T\}"



Clearly, "M_n^s(\R)" is the set of all symmetric matrix. we show that it is subspace of "M_n(\R)" .


Let, "a,b\in \R" and "A,B\in M_n^s(\R)" , thus "A^T=A,B^T=B"


Now, consider the liner combination "aA+bB" .


Note that


"(aA+bB)^T=(bB)^T+(aA)^T\\
=(aA)^T+(bB)^T\\
=aA^T+bB^T\\
=aA+bB"

Thus,we can see that


"aA+bB\in M_n^s(\R)"



Hence,it can be seen that



"M_n^s(\R)" is subspace of "M_n(\R)"